# Design and Evaluation of a PMT High-Voltage system for Deepsea Neutrino Telescope

### v1

@@ 3.2 PMT gain measurement and stability
Here, $\mathrm{Ped}(q)$ denotes the Gaussian electronic pedestal, $B$ is the normalization factor, and $\lambda$ is the mean detected photoelectron count. The SPE response function is parameterized following Ref. [DEAP:2017fgw] as a sum of two Gamma components and an exponential low-charge tail,
$\mathrm{SPE}(q)=\eta_{1}\,\Gamma(q;\mu,b)+\eta_{2}\,\Gamma(q;\mu f_{\mu},\,bf_{b})+\begin{cases}\eta_{3}\,\ell\,e^{-\ell q},&q<\mu,\\[2.0pt] 0,&q>\mu,\end{cases}$ (3.2)
where the first Gamma term describes full dynode multiplication [PRESCOTT1966173], the second captures reduced-charge events from incomplete multiplication at early dynodes, and the exponential term models the continuous low-charge population arising from elastic backscattering and other partial-amplification processes. This gamma family has been shown to reproduce the full SPE lineshape in many cases [DEAP:2017fgw, Akashi-Ronquest:2019mlk, Kalousis:2020PMTCalib].
A binned $\chi^{2}$ minimization was used to extract the ten free parameters governing the pedestal, SPE response, and Poisson occupancy. An example fitted charge spectrum is shown in Fig. 6. The gain uncertainty is obtained by propagating the parameter covariance from the SPE-spectrum fit to the mean SPE charge. This provides the statistical error on the extracted gain.
@@
$G(V)=\alpha V^{\beta},$ (3.3)
where $\beta$ characterizes the voltage sensitivity of the amplification. To determine the parameters $\alpha$ and $\beta$, the gain was measured at several high-voltage settings between 1100-1100 V and 1300-1300 V. The resulting gain–voltage relation is shown in Fig. 6. Because the gain–voltage response varies from PMT to PMT, measuring the gain curve for each unit allows the bias voltage to be tuned later to restore the target gain as the tube ages or its response drifts.
@@
After completing the base soldering procedure and assembling 31 PMTs into a single hDOM, the hDOM was continuously powered and operated for around 100 hours at a fixed high voltage (corresponding to a gain of $1\times 10^{7}$), in order to assess their long-term gain and baseline stability. Throughout this period, six data runs, each covering around 5 hours, were recorded using the motherboard designed for TRIDENT Phase-I [yangyong2025], which features a 125 MSps ADC. After determining the gain–voltage parameters, the bias of each PMT was set to achieve a uniform nominal gain. The long-term gain stability of the 31 PMTs in the hDOM was then evaluated using the dark-count data. Each recorded waveform spans 480 ns (60 ADC samples). The first run defines $t=0$ h, and the subsequent runs were taken at 9, 19, 43, 67, and 91 h. For each run, the plotted value represents the mean gain over the full 5-hour dataset, and the error bars indicate the the covariance of the spe gain fit.Fig. 7 shows the gain evolution over time. The gains stabilize after approximately 20 hours of continuous operation. For most channels, the variation remains within $\pm 2$$3\%$ of the initial value, and no systematic drift or monotonic trend is observed. Even the PMTs with the highest initial gains remain stable to within 5% throughout the measurement period. For the baseline stability of PMTs, the standard deviation of the first 20 samples was used to quantify the baseline fluctuation for each waveform. Fig. 8 shows the resulting stability over time. The run schedule is identical to that used for the gain measurement. For each run, the plotted value represents the mean baseline fluctuation over the full 5-hour dataset, and the error bars indicate the variation within that period.
The baseline stabilizes after roughly 40 hours of continuous operation, with channel-to-channel variations within 2%. This indicates negligible drift in the front-end electronic noise. Because baseline fluctuation directly affects the signal-to-noise ratio and effectively determines the single photoelectron (SPE) trigger threshold, these results demonstrate that the readout system maintains consistently low-noise performance suitable for long-duration deep-sea operations.
@@ 3.3 PMT Time Resolution
Precise photon timing is essential for neutrino telescopes, as the reconstruction of tracks and cascade directions and vertices relies on the arrival times of Cherenkov photons at optical modules. Simulation has shown that sub-degree angular resolution for muon tracks can be achieved only if the per-photon timing precision is on the order of 2 ns [Aguilar2011_ANTARES]. This requirement arises primarily from the intrinsic transit-time spread (TTS) of the PMT, the chromatic dispersion in seawater, and the undersea acoustic positioning accuracy. The PMT TTS, defined as the full width at half maximum (FWHM) of the distribution of single photoelectron (PE) transit times, is a key detector parameter to be evaluated.
@@
The TTS of each PMT is measured using the setup shown in Fig. 5. The laser intensity is adjusted to maintain a single-PE occupancy of approximately 0.05. PMT waveforms are digitized using the PicoScope sampling at 5 GSps. The laser system includes an internal optical splitter and a fast photodiode. An emitted laser pulse is divided into two paths: one eventually hits the PMT and the other hits the photodiode, which generates a reference electrical signal. The timing jitter of this reference channel, calibrated by the manufacturer, is less than 20 ps. This configuration allows sub-ns accuracy in the TTS measurement.
The photon arrival time on a PMT is defined as the point where the PMT waveform exceeds one third of SPE amplitude. For each event, the transit time (TT) is calculated as the time difference between the PMT signal and the reference laser trigger. A fixed offset caused by unequal propagation paths of the optical pulse and electrical reference signal was measured independently using a silicon photomultiplier (SiPM, SPTR $\sim$150 ps) [trident2024sipm] and is subtracted from PMT signals. A representative TTS measurement is shown in Fig. 9. The TTS is typically extracted either from the width of the Gaussian core or directly from the full distribution. As expected for small-mass PMTs, the distribution shows a mild asymmetry with a late-time tail arising from delayed electron trajectories, including elastic back-scattering at the first dynode.

### v2

@@ 3.2 PMT gain
where $\mathrm{Ped}(q)$ denotes the Gaussian electronic pedestal with two parameters $\mu_{\rm ped}$ and $\sigma_{\rm ped}$. $B$ is an overall normalization factor, and $\lambda\ll 1$ is the mean photoelectron occupancy.
The SPE charge response is modeled as a mixture of two Gamma components and an exponential low-charge tail [4]:
$\mathrm{SPE}(q)=\eta_{1}\,\Gamma(q;\mu,b)+\eta_{2}\,\Gamma(q;\mu f_{\mu},\,bf_{b})+\begin{cases}\eta_{3}\,\ell\,e^{-\ell q},&q<\mu,\\[2.0pt] 0,&q\geq\mu,\end{cases}$ (3.2)
where $\eta_{1,2,3}\geq 0$ are the component weights satisfying $\eta_{1}+\eta_{2}+\eta_{3}=1$. The first Gamma term describes the full dynode multiplication process and is characterized by the mean SPE charge $\mu$ and a shape parameter $b$ that controls the width of the full-multiplication peak [12]. The second Gamma term captures reduced-charge events from incomplete multiplication at early dynodes; its mean and shape are scaled relative to the primary component by $f_{\mu}$ and $f_{b}$, respectively. The exponential component, with decay constant $\ell$, models the continuous low-charge population arising from elastic backscattering and other partial-amplification processes. This Gamma-family parameterization has been shown to reproduce the full SPE lineshape in many cases [4, 3, 10].
The overall fit function (Eq. 3.1) is governed by a total of 10 free parameters: two pedestal parameters ($\mu_{\rm ped}$, $\sigma_{\rm ped}$), seven SPE-response parameters ($\mu$, $b$, $f_{\mu}$, $f_{b}$, $\eta_{2}$, $\eta_{3}$, $\ell$) with $\eta_{1}=1-\eta_{2}-\eta_{3}$ fixed by normalization, and one Poisson occupancy parameter ($\lambda$). A binned $\chi^{2}$ minimization was used to extract the ten free parameters governing the pedestal, SPE response, and Poisson occupancy. An example fitted charge spectrum is shown in Fig. 6 left. The gain uncertainty is obtained by propagating the parameter covariance from the SPE-spectrum fit to the mean SPE charge. This provides the statistical error on the extracted gain.
@@
$G(V)=\alpha V^{\beta}$ (3.3)
where $\beta$ characterizes the voltage sensitivity of the amplification. To determine the parameters $\alpha$ and $\beta$, the gain was measured at several high-voltage settings between 1100-1100 V and 1300-1300 V. The resulting gain–voltage relation is shown in Fig. 6 right. Because the gain–voltage response varies from PMT to PMT, measuring the gain curve for each unit allows the bias voltage to be tuned later to restore the target gain as the tube ages or its response drifts.
@@
3.3 Gain and Baseline Stability
The stability measurements were conducted using the integrated hDOM setup. Following the assembly of the 31 PMTs into a single hDOM, the hDOM was continuously powered and operated for around 100 hours at a fixed high voltage (corresponding to a gain of $1\times 10^{7}$) to assess their long-term gain and baseline stability. Throughout this period, six data runs, each covering around 5 hours, were recorded using the motherboard designed for TRIDENT Phase-I [17, 18]. The motherboard is controlled by a host PC via a high-speed optical fiber link. Both power and communication are delivered through an electro-optical cable connected to the hDOM via a specialized titanium feedthrough penetrator in the glass vessel.
After determining the gain–voltage parameters (Sec. 3.2), the bias of each PMT was set to achieve a uniform nominal gain. The long-term gain stability of the 31 integrated PMTs was then evaluated using dark-count data, with each recorded waveform spanning 480 ns (60 ADC samples). The initial run defines $t=0$ h, with subsequent measurements taken at 9, 19, 43, 67, and 91 h. Each data point in Fig. 7 represents the mean gain extracted from a 5-hour dataset, where the error bars denote the uncertainty derived from the SPE-fit covariance. The gains are observed to stabilize after approximately 20 hours of continuous operation. For most channels, the fluctuations remain within $\pm 2$$3\%$ of the initial values with no observable systematic drift. Even the channels with the highest gain settings maintain stability within 5% throughout the 100-hour test period.
The baseline stability analysis was conducted using the same dataset as the gain stability study. The standard deviation of the first 20 samples (160 ns) of each waveform was used to quantify noise fluctuations, with the error bars representing the standard deviation of these fluctuations within each run. This interval was selected from the pre-trigger buffer, a region preceding the arrival of the PMT pulse, ensuring that the estimate reflects the intrinsic electronic noise. As shown in Fig. 8, the baseline reaches a steady state after roughly 40 h of continuous operation, with channel-to-channel variations remaining within 2%. This indicates negligible drift in the front-end electronic noise once thermal equilibrium is established.
@@ 3.3 Gain and Baseline Stability
The Signal-to-Noise Ratio (SNR) was evaluated by comparing the SPE amplitude to the baseline noise. This ratio is critical for high detection efficiency and low false-trigger rates. Fig. 9 shows a representative SPE waveform at a gain of $10^{7}$. The pulse peak is $\sim 8.2$ mV against a baseline RMS of $\sim 0.2$ mV, yielding an SNR of $\sim 40$. This high signal fidelity ensures a stable trigger threshold, meeting TRIDENT requirements for long-term deep-sea operation.
Figure 9: Representative SPE waveform at $10^{7}$ gain recorded by the integrated hDOM setup.
@@ 3.4 PMT Time Resolution
Precise photon timing is essential for neutrino telescopes, as the reconstruction of tracks and cascade directions and vertices relies on the arrival times of Cherenkov photons at optical modules. Simulation has shown that sub-degree angular resolution for muon tracks can be achieved only if the per-photon timing precision is on the order of 2 ns [6]. This requirement arises primarily from the intrinsic transit-time spread (TTS) of the PMT, the chromatic dispersion in seawater, and the undersea acoustic positioning accuracy. The PMT TTS, defined as the full width at half maximum (FWHM) of the distribution of single photoelectron (PE) transit times, is a key detector parameter to be evaluated.
@@
The TTS of each PMT was measured using the laser test bench (Sec. 3.1) with the intensity adjusted to maintain a low SPE occupancy ($\lambda\approx 0.05$). PMT waveforms were digitized by a PicoScope at 5 GSps, providing 200 ps temporal binning. The timing jitter of the photodiode reference channel is less than 20 ps, as calibrated by the manufacturer. This configuration allows sub-ns accuracy in the TTS measurement.
The photon arrival time is defined as the point where the PMT waveform exceeds one-third of the averaged SPE peak amplitude. For each event, the transit time (TT) is calculated as the time difference between the PMT signal and the laser trigger. A fixed offset, arising from unequal optical and electrical propagation paths, was measured using a high-speed Silicon Photomultiplier (SiPM, SPTR $\sim$150 ps  [19]) and subtracted from the results.
A representative TTS measurement is shown in Fig. 10. The TTS is typically extracted either from the width of the Gaussian core or directly from the full distribution. As expected for small-mass PMTs, the distribution shows a mild asymmetry with a late-time tail arising from delayed electron trajectories, including elastic back-scattering at the first dynode.